OpenStudy (anonymous):

What number must you add to complete the square? x2 + 8x = -3

OpenStudy (anonymous):

(8/2)^2 = 16

OpenStudy (anonymous):

x^2 + 8x + 16 = -3 + 16 (x+4)^2 = 13